### v1

# PIPA: A Unified Evaluation Protocol for Diagnosing Interactive Planning Agents

Abstract
The growing capabilities of large language models (LLMs) in instruction-following and context-understanding lead to the era of agents with numerous applications. Among these, task planning agents have become especially prominent in realistic scenarios involving complex internal pipelines, such as context understanding, tool management, and response generation. However, existing benchmarks predominantly evaluate agent performance based on task completion as a proxy for overall effectiveness. We hypothesize that merely improving task completion is misaligned with maximizing user satisfaction, as users interact with the entire agentic process and not only the end result. To address this gap, we propose PIPA, a unified evaluation protocol that conceptualizes the behavioral process of interactive task planning agents within a partially observable Markov Decision Process (POMDP) paradigm. The proposed protocol offers a comprehensive assessment of agent performance through a set of atomic evaluation criteria, allowing researchers and practitioners to diagnose specific strengths and weaknesses within the agent’s decision-making pipeline. Our analyses show that agents excel in different behavioral stages, with user satisfaction shaped by both outcomes and intermediate behaviors. We also highlight future directions, including systems that leverage multiple agents and the limitations of user simulators in task planning.
1 Introduction
Large language models (LLMs) are increasingly deployed in real-world applications, primarily due to their ability to understand complex goals and devise structured sequences of action: a process known as “planning” (Wang et al., 2024). To assess and refine planning skills, researchers have introduced diverse benchmarks in web-based (Yao et al., 2022; Xie et al., 2024b), mobile (Deng et al., 2024), embodied (Choi et al., 2024), and automated testbeds (Park et al., 2023).
One of the most impactful planning applications is task planning, where agents generate and execute domain-specific plans (e.g., itineraries) to help users achieve goals while adhering to contextual constraints (Budzianowski et al., 2018; Rastogi et al., 2020; Xie et al., 2024a). While agents must accurately interpret user requests, leverage predefined tools, and engage in personalized dialogues, many benchmarks focus solely on final task completion, overlooking the agents’ intermediate behaviors and planning steps. Given that overall user satisfaction is shaped by multiple factors (e.g., efficiency and effectiveness) throughout prolonged interactions (Ashfaq et al., 2020), a narrow focus on outcome-based metrics can misrepresent an agent’s true effectiveness, especially in complex task-oriented applications (Cao et al., 2025; Xia et al., 2025; Acikgoz et al., 2025).
In addition, existing benchmarks across different domains often introduce bespoke evaluation frameworks and metrics, making it possible for the same agent to be evaluated using entirely different criteria. This fragmentation introduces inconsistencies in how agent behavior is interpreted and makes it difficult to draw generalizable conclusions about agent capabilities. Without a unified evaluation protocol that captures key aspects of agent behavior beyond task-specific outputs, it becomes challenging to meaningfully compare systems, track progress over time, and identify fundamental limitations in agent design.
To this end, we propose PIPA, a unified evaluation Protocol for Interactive Planning Agents. To generalize the evaluation criteria across arbitrary agentic scenarios, PIPA is formulated within the framework of a partially observed Markov Decision Process (POMDP), recognizing that real-world states (e.g., user intentions and satisfactions) are often hidden or only partially observable. By capturing these dynamics, PIPA defines domain- and task-agnostic evaluation criteria that can be easily instantiated through a set of atomic and easily measurable metrics. As illustrated in Figure 1, PIPA provides the following three key advantages:
  1. 1.
    Unified Evaluation: PIPA establishes a task-agnostic protocol for interactive task planning, enabling a consistent comparison across diverse benchmarks under a shared set of principles, which has been discussed for decades (Walker et al., 1997) but still has not been fully resolved.
  2. 2.
    Multi-Axis Diagnosis: By aligning with the POMDP framework, PIPA supports evaluation along multiple stages of the interaction pipeline, encompassing both intermediate decisions and final outcomes. This allows for a fine-grained diagnostic analysis of how different agents behave and impact performance.
  3. 3.
    Cross-Benchmark Comparisons: PIPA enables systematic comparisons across heterogeneous tasks and environments, thereby uncovering planning strategies and revealing performance trade-offs that may not be observable within the scope of individual benchmarks.
Figure 1: PIPA provides domain- and task-agnostic diagnostic evaluation protocols for interactive task agents following the POMDP paradigm.
Through extensive experiments, we demonstrate that different models exhibit distinct strengths and weaknesses within decision-making pipelines. Furthermore, human studies indicate that stage-specific evaluation using PIPA correlates more strongly with improvements in user satisfaction than with the final task completion metric alone. Lastly, analyses on combining agents during deployment and the reliability of user simulators highlight promising research directions for future scholars and practitioners.
2 Background: Bespoke Evaluation of Planning Agents
Recent benchmarks for evaluating agentic tasks highlight a variety of metrics tailored to capture different facets of agent performance, reflecting the complexity and multi-dimensional nature of interactive planning tasks, as illustrated in Table 5 in Appendix A. For instance, AgentBench (Liu et al., 2024) incorporates domain-specific metrics such as Success Rate, Win Rate, F1 Score, and Exact Match to assess agents comprehensively. Similarly, WebLINX (Lù et al., 2024) introduces granular evaluation dimensions, including Intent Match, a boolean indicator of correct intent prediction; Element Similarity, verifying correct function arguments; and Text Similarity, focusing on the lexical consistency of arguments. FlowBench (Xiao et al., 2024) further enhances the evaluation landscape by adding metrics focused explicitly on tool usage: Tool Invocation, measured by precision, recall, and F1 score for identifying the correct tool configurations; Success Rate, denoting the proportion of entirely successful sessions; and Task Progress, capturing the percentage of goals completed within a session. Meanwhile, $\tau$-Bench (Yao et al., 2025) utilizes more direct measures of success with Pass@k, a boolean indicator reflecting success across k attempts, and Pass^k, which requires successful outcomes for all k trials. Unlike PIPA, all such works tailor their evaluation metrics to their usage, thus lacking in a generalizable evaluation across benchmarks. Furthermore, the primary focus of most benchmarks is measuring task completion rates, which has been pointed out for several years (Ribeiro et al., 2020). Even for complex and subjective tasks such as information aggregation, the evaluation is done with final task success (Reddy et al., 2024; Yoran et al., 2024).
3 Proposed Evaluation Protocol: PIPA
As a prerequisite step, we set minimum requirements for task planning benchmarks to effectively show realistic scenarios. Specifically, task planning benchmarks should define specific tools (e.g., APIs) and construct domain-specific databases. These assumptions are adopted because their dedicated environmental constraints present challenges in achieving a unified evaluation.
3.1 Components of PIPA
Metric Design Criteria
Figure 2: Decision theory-based taxonomy of evaluation metrics in PIPA. As mentioned in Walker et al. (1997), it should be cautious to generalize the original taxonomy to different agents and tasks. Following this, we provide distinct interpretations of each element, which will be described in further sections.
Interactive planning tasks are generally modeled using the POMDP paradigm, where agents devise plans based on partial observational data (Sun et al., 2024; He et al., 2024; Xie et al., 2024a; Yao et al., 2025). Building on the framework of Williams & Young (2007), we present an abstract pipeline consisting of $\mathcal{S}$ (a set of agent states), $\mathcal{A}$ (a set of actions), $\mathcal{O}$ (a set of observations) and $\mathcal{R}$ (a set of rewards) [^1] Our formulation adopts discrete spaces while acknowledging the existence of continuous concepts, such as transition and observation probabilities., along with an optional policy definition $\mathcal{P}$ (a set of global policies). The metrics of PIPA are defined for each phase within this pipeline, as various factors contribute to user satisfaction beyond a singular overall evaluation (Ashfaq et al., 2020; Siro et al., 2022). Consequently, PIPA aims to provide a comprehensive evaluation of agent performance, not just in terms of the overall results but also by diagnosing which specific phases may need improvement. Considering the fact that most behavioral stages of LLM agents are pipelined with natural languages, we employ a set of atomic LLM-based evaluations by following previous observations where atomic and decomposed evaluation criteria yield consistent and reliable outcomes (Liu et al., 2023; Saad-Falcon et al., 2024; Lee et al., 2025).
Furthermore, we align PIPA with the decision theory-based taxonomy (Walker et al., 1997), as illustrated in Figure 2. In line with this taxonomy, the evaluation metrics in PIPA (i.e., state consistency, tool efficiency, observation alignment, policy adherence, and task completion) contain a “disparate set of performance measures” that can collectively assess the fundamental objective: maximizing user satisfaction. By establishing a principled connection between decision theory and POMDP-based PIPA, we ground our evaluation framework in a principled model of sequential decision making under uncertainty. Specifically, we leverage the fact that POMDPs are a natural instantiation of decision theory, where the agent selects actions to maximize expected utility given a belief over latent states. This integration enables PIPA to measure user satisfaction holistically while remaining grounded in theoretical rigor.
Throughout this work, as visualized in Figure 1, we formulate the basic components of multi-turn, multi-step agents as follows: $T$ denotes the total number of turns (i.e., alternating utterances between the user and the agent) in an interactive session, $M_{t}$ represents the number of steps in the agent’s internal reasoning process within a single utterance at turn $t$, and $\{U_{\tau},A_{\tau}\}_{\tau=1}^{t}$ signifies the sequence of user inputs and agent responses observed up to turn $t$.
3.1.1 State Consistency ($\mathcal{S}$)
The intermediate step of the agent is pivotal in summarizing user requests and determining optimal subsequent actions (Shin et al., 2022; Wei et al., 2022). Given that both pipelined and end-to-end multi-turn interactions are inherently susceptible to error accumulation (Liao et al., 2018; 2021), it is crucial to validate whether intermediate outcomes consistently mediate between user inputs and agent outputs. In this context, the state consistency metric measures whether an agent correctly aligns user requests with its intermediate steps ($S_{t}$). Notably, agent’s intermediate steps can be represented in either a “structured format” (e.g., dialogue states) or “natural language” (e.g., Chain-of-Thought). By accommodating both representations, the state consistency can be formulated as follows:
$\text{{PIPA}}_{\mathcal{S}}=\frac{\displaystyle\sum_{t=1}^{T}\sum_{k=1}^{M_{t}}\text{IsConsistent}\bigl{(}S_{t}^{k},\;[\{U_{\tau},A_{\tau}\}_{\tau=1}^{t-1};U_{t}]\bigr{)}}{\displaystyle\sum_{t=1}^{T}M_{t}}$ (1)
Here, $S_{t}^{k}$ denotes the $k$-th state within the turn $t$. In addition, the IsConsistent function serves as a boolean indicator evaluated by LLMs that compares user requests (mostly natural language) with internal states (structured or natural language) [^2] Practically, as observed by Lee et al. (2025), evaluating a group of targeted samples collectively, using simple and atomic criteria such as boolean assessments, does not adversely affect performance compared to stepwise evaluation. Therefore, we group $S_{t}^{1:k}$ in the implementation..
3.1.2 Tool Efficiency ($\mathcal{A}$)
The management of external tools or functions, typically in the form of APIs, incurs operational costs and affects the task performance, underscoring the importance of evaluating their effective utilization. In particular, the occurrence of failed API calls prior to collecting complete information from users–commonly observed in multi-turn interactions–can result in unnecessary expenditures of resources (i.e., time and money) for service providers. To address this, the tool efficiency metric includes the total number of tool calls ($C_{T}$). For a given task, agent scenarios that require fewer API calls are considered indicative of more efficient API management.
Additionally, due to the stochastic nature of natural language prompting, generating well-structured API calls poses a significant challenge for certain agents. Even when an agent invokes an API at an appropriate time, iterative attempts at call generation will eventually lead to additional, avoidable costs. Consequently, this metric also calculates the number of failed tool generation attempts ($C_{F}$). Formally, we propose the following tool efficiency metric:
$\text{{PIPA}}_{\mathcal{A}}=\frac{C_{T}-C_{F}}{C_{T}+C_{F}}$ (2)
where the numerator rewards successful calls and the denominator penalizes excessive or failed calls. A higher value of $\text{{PIPA}}_{\mathcal{A}}$ indicates more efficient tool usage: maximizing successful calls while minimizing total and failed calls.
3.1.3 Observation Alignment ($\mathcal{O}$)
According to the original categorization of “qualitative measures” in Figure 2, criteria such as inappropriate response were used to indicate when the system’s output was factually, contextually, or semantically at odds with user’s needs (Walker et al., 1997). While such qualitative assessments provide valuable interpretative insights, they are difficult to incorporate directly into a formal decision‐theoretic framework. Consequently, we transition from a purely qualitative to a more quantitative representation so that the measure can be more easily integrated with our reward-based evaluation [^3] Crucially, this shift does not abandon the interpretive depth of the original qualitative measures; rather, it refines them into a metric that preserves underlying meanings and enables consistent, systematic evaluation across different dialogues and use cases..
On top of this discourse, the observation alignment metric evaluates whether observations appearing within the agent responses align with what the user requires. Specifically, it is calculated with a boolean criteria for each observation (i.e., retrieved database entity), preceded by capturing the number of observations ($|O|$) explicitly appear within the agent responses ($R_{agent}$). Considering the fact that both user utterance and agent response are represented in natural languages, it can be measured via a set of atomic LLM evaluations:
$\text{{PIPA}}_{\mathcal{O}}=\frac{1}{|T_{\text{obs}}|}\sum_{t\in T_{\text{obs}}}\left(\frac{1}{|O^{(t)}|}\sum_{o\in O^{(t)}}\text{IsAligned}(o,[\{U_{\tau},A_{\tau}\}_{\tau=1}^{t-1};U_{t}])\right)\,,$
$\text{where}\;T_{obs}=\{t\in T:|O^{(t)}|>0\}$ (3)
Intuitively, $T_{obs}$ denotes a set of turns where observations are present within agent responses ($|O|>0$), and observations of a specific turn $t\in T_{obs}$ are extracted from LLM (denoted $O^{(t)}$). In addition, IsAligned serves as a boolean indicator assessing whether each observation aligns with conversational context. By quantifying and improving observation alignment, we not only promote the clarity of agent outputs but also minimize the system’s overall costs by decreasing the possibility of repetitively calling tools.
3.1.4 Optional: Policy Alignment ($\mathcal{P}$)
Some interactive benchmarks assume certain policies (a behavioral rule appearing in agent responses, not only emerging in their internal states) that are globally reflected across the interactive sessions, such as Xiao et al. (2024). While some agents fail to provide tasks due to non-compliance with these policies, there are cases where agents do not follow the policy but succeed in achieving the intended goal (e.g., the policy “not recommending more than three items at a time” does not directly affect the task performance). Policy alignment is a session-level metric measuring whether a predefined set of policies ($\mathcal{P}$) are consistently followed throughout interactive sessions.
$\text{{PIPA}}_{\mathcal{P}}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}{\text{IsAdherent}(p;\{U_{t},A_{t}\}_{t=1}^{T})}$ (4)
Similar to the state consistency metric (Section 3.1.1), IsAdherent serves as a boolean indicator deciding consistency between each policy and interaction context.
3.1.5 Task Completion ($\mathcal{R}$)
The primary objective of a task agent is to effectively accomplish goal-oriented tasks. The task completion metric adheres to the task-specific performance criteria established by each benchmark, since benchmark scenarios typically define their own evaluation frameworks to determine what constitutes a “completed task.” Specifically, most metrics in diverse benchmarks demonstrated in Table 5 can be regarded as task completion metrics since they mainly focus on whether agents successfully accomplish the given task.
3.1.6 Agent Interaction Pattern
While prior work has typically reported on dataset-level statistics (e.g., basic information in Table 1), it has rarely addressed how agents actually interact within their environments. We suggest reporting the following information, providing an overview of the interactive tendency of employed agents in their specific scenarios.
The Number of Turns
Different from offline dialogue datasets where a fixed number of turns is provided, recently employed online evaluation scenarios typically omit details about the agent interactivity. However, the number of interactive turns serves as a critical indicator of the conversational or interactive dynamics. In general, planning scenarios with more turns tend to require further management of user constraints, suggesting the need for sustained engagement and state tracking.
The Number of Steps
Steps refer to internal actions or reasoning hops, such as goal decomposition or tool use. This feature reflects the agent’s ability to handle complex tasks that require breaking down overarching goals into smaller, manageable actions. Fewer steps may suggest shallow or overly generic reasoning, while more steps indicate finer control. Correlating the count of steps with performance and the length of interaction can reveal whether agents are making meaningful progress or getting stuck in unproductive loops.
4 Experiments
4.1 Benchmarks
# of Dialogues # of Tools # of Database
TravelPlanner (2024a) 180 7 3,865,195 total, 3,827,361 max for a tool
$\tau$-Bench-Airline (2025) 50 13 500 users, 300 flights, 2,000 reservations
$\tau$-Bench-Retail (2025) 115 15 500 users, 50 products, 1,000 orders
Table 1: Benchmark statistics. We use a validation set for TravelPlanner that demonstrates a similar performance pattern with the test set.
To verify the effectiveness of the proposed evaluation protocol, we employ challenging task planning benchmarks that conform to the prerequisite assumptions described in Section 3 (i.e., predefined tools and external database). Specifically, we utilize three task domains in two benchmarks: TravelPlanner (Xie et al., 2024a), a single-turn dataset in the itinerary domain, and $\tau$-Bench (Yao et al., 2025), a multi-turn dataset encompassing the airline and retail domains. Table 1 presents the statistics of these benchmarks. For the evaluation metric for task completion ($\mathcal{R}$), pass rate is used for TravelPlanner, and pass^k is used in $\tau$-Bench. Refer to Table 5 for the description of each metric.
4.2 LLMs for Agent, User Simulator, and Performance Evaluator
TravelPlanner $\tau$-Bench-Airline $\tau$-Bench-Retail
[…]
Proprietary Models
gpt-4o .98 .79 .01 .48 .01 .44 .58 .96 .73 .83 .42 .70 .53 .93 .79 .89 .54 .74
gpt-4o-mini .94 .95 .00 .50 .00 .48 .49 .92 .71 .84 .26 .64 .48 .89 .72 .89 .46 .69
gpt-3.5-turbo .95 .93 .00 .50 .00 .48 .56 .79 .65 .73 .08 .56 .44 .69 .57 .78 .21 .54
gemini-1.5-fsh. .99 .71 .00 .31 .00 .40 .55 .84 .66 .87 .32 .65 .60 .85 .72 .86 .15 .64
sonnet-3.5 .99 .99 .01 .56 .01 .51 .80 .97 .84 .87 .46 .79 .73 .94 .85 .91 .62 .81
Open-Source Models
mistral-large .87 .77 .00 .36 .00 .40 .31 .96 .50 .76 .26 .56 .23 .91 .44 .80 .34 .54
mixtral-8x7B .94 .77 .00 .27 .00 .40 .25 .85 .17 .64 .28 .44 .27 .50 .27 .65 .05 .35
llama-3.3-70B .95 .96 .00 .39 .00 .46 .38 .90 .70 .81 .30 .62 .32 .88 .79 .87 .36 .64
qwen2.5-72B .98 .97 .00 .39 .00 .47 .35 .90 .58 .79 .20 .56 .27 .75 .53 .84 .38 .55
Table 2: PIPA evaluation result. The relative performance ranking among agents for each metric is differentiated with colors (i.e., darker color indicates more competitive performance; different colors are applied for proprietary and open-source models). Best average performance is indicated in bold, and second-best is underlined.
TravelPlanner $\tau$-Bench-Airline $\tau$-Bench-Retail
Avg. Turns Avg. Steps Avg. Turns Avg. Steps Avg. Turns Avg. Steps
Proprietary Models
gpt-4o 1.00±0.00 18.71±6.91 7.70±3.08 0.93±1.71 8.26±2.54 0.98±1.23
gpt-4o-mini 1.00±0.00 15.16±6.67 8.72±3.77 0.81±1.79 8.52±3.09 1.07±1.50
gpt-3.5-turbo 1.00±0.00 12.87±6.44 7.66±3.70 0.94±1.12 7.97±2.50 1.17±1.19
gemini-1.5-fsh. 1.00±0.00 16.57±6.97 3.04±2.14 0.47±0.70 6.74±4.77 0.44±0.68
sonnet-3.5 1.00±0.00 18.36±6.84 6.20±2.10 1.03±1.58 7.71±2.06 1.15±1.49
Open-Source Models
mistral-large 1.00±0.00 14.16±7.48 8.46±6.33 0.62±1.74 12.51±6.44 0.68±0.96
mixtral-8x7B 1.00±0.00 18.44±7.84 18.14±9.99 0.01±0.12 16.11±9.69 0.01±0.09
llama-3.3-70B 1.00±0.00 18.52±6.54 8.02±4.33 1.22±1.63 6.44±4.03 1.31±1.31
qwen2.5-72B 1.00±0.00 18.20±7.36 18.74±9.21 0.04±0.86 17.10±9.73 0.01±0.09
Table 3: The average number of turns and steps within a single interactive session (the agent interaction pattern defined in Section 3.1.6). Note that Avg. Steps denotes the number of internal agentic processes in each turn not visible in superficial interactions.
We tested diverse task planning agents consisting of five proprietary models (gpt-4o, gpt-4o-mini, gpt-3.5-turbo, gemini-1.5-flash, sonnet-3.5) and four open-source models (mistral-large 444123B-sized model: https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, mixtral-8x7B, llama-3.3-70B, qwen2.5-72B). Since $\tau$-Bench requires multi-turn interactions between a user and agent, we employ gpt-4o as a user simulator with the same instruction as the original work (Yao et al., 2025). For the LLM evaluator leveraged in measuring PIPA metrics, we employ llama-3.3-70B (Grattafiori et al., 2024) [^5] We confirm llama-3.3-70B shows better performance compared to gpt-4o-mini through our preliminary investigation.. All prompts used in our experiments are listed in Appendix B.
4.3 Results
Behavioral Stage Diagnosis
Table 2 presents the evaluation results following PIPA  [^6] We report agents’ erroneous behaviors at each stage in Appendix C.. We differentiate the color ranking of proprietary and open-source models to observe diverse patterns. Along with individual metrics in PIPA, we report the average score to see the overall behavioral performance.
A notable observation is that the conventional task completion metric ($\mathcal{R}$) does not necessarily reflect the performance of intermediate phases. For instance, while qwen2.5-72B achieves the highest performance on task completion (.38) among open-source models in $\tau$-Bench-Retail, its PIPA metric results lag behind other models. Moreover, although most agents in TravelPlanner show the same zero task completion performance, each model has different intermediate capability patterns according to PIPA metrics. These observed patterns imply a potential discrepancy between task completion performance values and qualitative user preferences, as user satisfaction is influenced not only by task completion but also by the quality of interactions across all phases. We introduce a human study for verifying this discrepancy in Section 5.1.
Interaction Pattern Diagnosis
We summarize each agent’s interaction pattern in Table 3. While TravelPlanner is characterized by single-turn interactions with numerous intermediate steps prior to producing a final response, $\tau$-Bench exhibits fewer intermediate steps per turn but involves multiple turns. These patterns reveal certain behavioral tendencies among the agents; for instance, competitive agents (e.g., sonnet-3.5, llama-3.3-70B) appear to engage in more extensive internal reasoning within individual turns, yet participate in fewer turns overall in multi-turn benchmarks. This observation suggests that, in adapting TravelPlanner for multi-turn scenarios, careful consideration must be given to balancing the number of turns and steps within each turn. Given that service providers often target different application scenarios, analyzing interaction patterns across benchmarks can inform the strategic selection of those most aligned with specific deployment contexts.
5 Analyses and Discussions
5.1 Relationship Between Task Completion
and User Satisfaction
Our hypothesis regarding the results presented in Table 2 is that factors beyond task completion may have significantly influenced overall user satisfaction. To investigate this, we conduct a human study involving 16 graduate-level participants across the following two controlled scenarios:
(1) Same $\mathcal{R}$ Different AVG.
We select gpt-4o-mini and mistral-large in $\tau$-Bench-Airline, as these two models exhibit identical task completion performance (.26), yet differ in their PIPA average scores (.64 and .56, respectively).
(2) Same AVG. & Different $\mathcal{R}$
We select gemini-1.5-fsh. and llama-3.3-70B in $\tau$-Bench-Retail, as these two models exhibit identical PIPA average score (.64), yet differ in their task completion performance (.15 and .36, respectively).
Participants are presented with two multi-turn conversations generated by different agents interacting with a gpt-4o user simulator. They are asked to indicate which conversation they find more satisfactory [^7] Detailed instructions are provided in Section D.1, and we conduct qualitative analysis for participants’ decisions in Section D.2.. If both conversations are perceived as equally (un)satisfactory, participants are given the option to select “no preference.” Each participant evaluates a total of six conversation pairs (randomly sampled three pairs of conversations, respectively).
Figure 3: Human study results for measuring preferred interactions.
The results of the human evaluation are presented in Figure 3. For both scenarios, we conduct binomial tests by excluding “no preference” responses, yielding statistically significant results (p-value $<$ 0.05 in both cases). In the Airline domain study, we observe that participants tend to favor conversations with higher average score when task completion is held constant. Conversely, in the Retail domain study, task completion still remains a strong predictor of user satisfaction when the average score is identical. These findings highlight the complementary roles of mostly-used task completion and PIPA metrics in capturing different facets of user satisfaction, as demonstrated in Table 2. Our study thus supports the use of PIPA as a meaningful proxy for human preferences and motivates future research into evaluation frameworks that go beyond simple task outcomes.
5.2 Can Mixing Agents Lead to Better
Performance?
Building upon the insights from previous experiments, we are further motivated to explore the potential for enhancing performance and user satisfaction by leveraging diverse agents strengths within the agentic pipeline. To facilitate this investigation, we separate the agent’s interactive process into two distinct components: (1) intermediate understanding and (2) response generation. This separation enables targeted improvements, wherein state consistency and tool efficiency are addressed in the first stage, whereas observation alignment and policy alignment are the focus of the second. Specifically, we employ one agent to understand user requests and call proper tools, then replace it with the other when starting to generate responses.
[…]
mistral-large .383±0.038 .924±0.046 .581±0.022 .777±0.005 .358±0.029 .605
   + qwen2.5-72B interm. .417±0.031 .939±0.017 .545±0.029 .809±0.008 .336±0.050 .609
qwen2.5-72B .403±0.021 .951±0.015 .637±0.031 .821±0.008 .329±0.021 .628
   + llama-3.3-70B interm. .447±0.020 .958±0.014 .663±0.022 .831±0.008 .272±0.051 .634
   + mistral-large interm. .423±0.039 .933±0.032 .578±0.050 .776±0.009 .367±0.013 .615
llama-3.3-70B .463±0.055 .952±0.013 .643±0.035 .818±0.015 .263±0.056 .628
   + qwen2.5-72B interm. .404±0.033 .953±0.007 .599±0.016 .803±0.014 .361±0.019 .624
Table 4: Selected results for mixing agent experiments in $\tau$-Bench-Airline. Agents indicated with “interm.” are utilized in an intermediate understanding component. Full results and additional discussions are presented in Table 7 and Section 5.2.
For faster investigation of this experimental study, we utilize FP8-quantized versions of three open-source models (llama-3.3-70B, qwen2.5-72B, mistral-large), alongside llama-3.3-70B as the user simulator, and evaluate on $\tau$-Bench-Airline with five repeated runs per configuration. Selected results are presented in Table 4. While acknowledging that arbitrary combinations of agents do not consistently yield improved performance, even leading to drop in performance, certain pairings demonstrate notable enhancements in either task completion or the average PIPA score. Notably, combining qwen2.5-72B (for intermediate understanding) with llama-3.3-70B (for response generation) leads to a substantial improvement in task completion (.263 $\rightarrow$ .361) without compromising the average PIPA score too much (.628 $\rightarrow$ .624). These findings suggest a promising direction for future research to strategically arrange agents across different stages of the task planning pipeline. For additional details and comprehensive experimental results beyond the scope of this discussion, please refer to Appendix E.
5.3 Analysis on the Reliability of User Simulator
To highlight potential limitations of multi-turn evaluation tasks, including those in our own study, we conduct a thorough analysis of user-agent conversations from $\tau$-Bench-Airline, where gpt-4o served as the user simulator. We manually examine all user utterances against corresponding user instructions to assess whether the user’s behavior aligns with the intended instructions.
Our analysis reveals that in 11 out of 50 conversations (22%), the user simulator demonstrates behaviors not in line with its instructions. While these deviations do not necessarily lead to failures in task completion, they highlight the importance of considering user simulator performance when evaluating an agent. Although addressing this limitation is left for future work, we identify the following erroneous patterns where a user simulator does not adhere to its instructions: (1) proactivity, (2) instruction contradiction, (3) missing details, and (4) misinterpretation. A comprehensive description of these error categories is provided in Appendix F.
6 Conclusion
We introduce PIPA, a domain- and task-agnostic evaluation protocol grounded in a partially observable Markov Decision Process (POMDP) framework, to capture how user satisfaction is shaped by both final outputs and intermediate behaviors in interactive task planning. Our experiments reveal that different models excel at various stages, underscoring the need for a holistic view of agent performance. Human studies also support that focusing solely on task completion can be misleading; stage-specific indicators often correlate more strongly with the user’s overall experience. These findings demonstrate that PIPA offers a principled, diagnostic lens to help researchers and practitioners design and benchmark more robust, transparent, and user-centric planning agents.
Appendix A Evaluation Metrics Across Task Planning Benchmarks
Benchmark Metric Description
Multi-Turn Planning Benchmarks
MultiWOZ (2018) Success Rate The system answered all requested attributes.
Inform Rate The system has provided an appropriate entity.
AgentBench (2024) Domain-Specific Different metrics such as Success Rate, Win Rate,
F1 Score, Exact Match, etc. are adopted.
WebLINX (2024) Intent Match Boolean indicator for correct intent prediction.
Element Similarity Correctly predicted function arguments.
Text Similarity Lexical similarity of arguments across functions.
FlowBench (2024) Tool Invocation Correctly identified tool configs (P/R/F1).
Success Rate Proportion of completely successful sessions.
Task Progress Percentage of completed goals within a session.
$\tau$-Bench (2025) Pass@k Boolean success indicator for k attempts.
Pass^k Whether all k trials are successful.
Single-Turn Planning Benchmarks
WebShop (2022) Task Score The average reward across episodes.
Success Rate The proportion of fully rewarded instructions.
TravelPlanner (2024a) Delivery Rate The agent completed tasks within limited steps.
Pass Rate The agent satisfied all plans and constraints.
Table 5: Performance evaluation metrics of service planning benchmarks. Note that metrics sharing the same title (e.g., Success Rate) have distinct definitions across different benchmarks.
Appendix B Prompts
Basically, we keep the same agent prompt as the original paper (Xie et al., 2024a; Yao et al., 2025). The following prompts are evaluation prompts adopted in PIPA evaluation.
LLM Evaluation Prompt for State Consistency
Instruction: You are tasked with evaluating whether each agent’s intermediate state (which can be either a thought-the agent’s internal reasoning-or an action-an API call) accurately reflects and mediates between:
1. The user’s requests (in the dialogue so far).
2. Any previously established agent states.
Your evaluation should focus on whether the agent’s intermediate steps exhibit clear, consistent reasoning that aligns the user’s inputs with the agent’s outputs, without introducing errors or contradicting earlier information.
Evaluation Criteria:
1. Consistency with the user request
- Does this state correctly respond to or reflect the user’s specific request(s) in the dialogue?
- Does the thought or chosen action remain faithful to what the user asked for?
2. Consistency with previous states
- Does this state align with earlier states (both thoughts and actions) without contradicting or omitting essential information?
- Does the progression of reasoning or actions flow logically from prior context?
3. Accuracy and truthfulness
- Does the state maintain factual correctness, avoiding hallucinations or irrelevant information?
- Does it accurately represent any data or entities referenced so far?
Scoring:
- 1 if the intermediate state is entirely consistent and correct (no contradictions, omissions, or factual errors).
- 0 if the state demonstrates any errors, contradictory information, missing critical details, or misalignment with the user’s request or prior states.
Dialogue History:
A chronological sequence of user and agent messages in JSON format, each with a "role" and "content."
{dial_history}
Agent’s States:
The agent’s states in JSON format, in chronological order after the dialogue history. Each state has a state_id, type, and content.
{states}
Output Format (JSON):
Return a list of objects in JSON, each containing:
"state_id": "1",
"justification": "brief explanation...",
"score": "0"
"state_id": "2",
"justification": "brief explanation...",
"score": "1"
LLM Evaluation Prompt for Observation Alignment
Instruction:
You are tasked with evaluating whether agent’s each response accurately aligns with the prior conversational history and other details if any. Specifically, you must verify:
- That any observations or entities referred to in the agent’s response meaningfully match the user’s stated requirements.
- That the number or scope of these observations is appropriate for the request (keeping in mind that varying amounts of recommendations or offerings can still be valid).
- That no contradictory, extraneous, or irrelevant observations are introduced.
Evaluation Criteria:
1. Consistency with the user request
- Do the observations (e.g., recommended items or database entities) and their details align with the user’s explicit request or needs?
- Are references to these observations relevant, or do they drift from the user’s stated goals?
2. Completeness relative to the request
- Are all key observations needed to fulfill the user’s request addressed, without omission of crucial details?
- If fewer (or more) observations are presented, is the choice justifiable in context?
3. Accuracy and truthfulness
- Are the observations factual, given the user’s query and available context?
- Does the response avoid hallucinated or incorrect data?
4. Consistency with previous details
- Does each current agent response remain consistent with all previously established facts or user-provided details?
- Are there no contradictions or misrepresentations of earlier statements?
Scoring:
- 1 if the agent’s response is fully consistent, addresses the request, and properly references any relevant observations (no errors or omissions).
- 0 if the response includes incorrect, missing, or misaligned observations, introduces contradictions, or strays from the user’s request.
Dialogue History:
Chronological user-agent messages in JSON format, each with a role and content.
{dial_history}
Agent’s Response:
The agent’s responses in JSON format, in chronological order. Each response has a response_id and content.
{responses}
Output Format (JSON):
Return a list of JSON objects, each containing:
"response_id": "1",
"justification": "brief explanation...",
"score": "0"
"response_id": "2",
"justification": "brief explanation...",
"score": "1"
LLM Evaluation Prompt for Policy Alignment
Instruction:
You are evaluating whether an agent consistently follows a predefined set of policies throughout the interactive session. Your primary task is to check the dialogue for any point where the agent might violate a policy.
In your Justification, mention the policy and turn message, if violated. Otherwise, mention all policies are adhered to.
1. If the agent violates any policy (even once), return Score: 0.
2. If the agent fully complies with all policies during the session, return Score: 1.
Dialogue History:
Chronological user-agent messages in JSON format, each with a role and content.
{dial_history}
Policy:
{policy}
Output Format (JSON):
Return a list of JSON objects, each containing:
"policy_id": "1",
"justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.",
"score": "0"
"policy_id": "2",
"justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.",
"score": "1"
Appendix C Qualitative Examples of Agents’ Erroneous Behaviors
Failure in State Consistency
USER: I want to know how many t-shirt options are available in in the store right now. AGENT STATES: - Thought: The user wants to know the number of t-shirt options available in the store. I can use the list_all_product_types tool to get the list of all product types, and then count the number of t-shirt options. - Action: get_order_details(order_id=#W0000000) Explanation: The agent demonstrated a failure in state consistency through how they made an action to get the user’s order details, when they should have provided user information about what was available in store.
Failure in Tool Efficiency
AGENT: Can you please provide me with your user ID? USER: My user ID is mia_li_3668 TOOL CALL: book_rservation(user=mia_li_3668, payment_id: credit_card_123456, ... ) TOOL RESPONSE: Error: payment method credit_card_123456 not found AGENT: Can you please confirm the payment method you would like to use? USER: I would like to use my two certificates to pay Explanation: The agent prematurely made tool calls before gathering the required information. In this example, it guesses a value for user id and payment id, before gathering this information from the user.
Failure in Observation Alignment
AGENT: Which payment method would you like to use for processing the refund? USER: You can use the certificate for processing the refund. AGENT: The refund has been processed to your gift card! Explanation: We observe a failure in observation alignment because the user specified to process the refund using the certificate, but the agent response states the refund was processed to the gift card.
Failure in Policy Adherence
Agent Policy 1: The agent must first obtain the user ID, then ask for the trip type, origin, and destination. USER: Im looking to book a flight from New York to Seattle on May 20 AGENT: I see there are three airports in the New York area... which do you prefer to depart from? The destination airport will be SEA. Explanation: The agent fails to obtain the user id before asking for the trip type, origin, and destination, as specified by the policy.
Table 6: Excerpt of failure cases in each PIPA metric.
Appendix D Details on Human Study
D.1 Instruction Given to Participants
Figure 4: A guideline provided to human participants in Section 5.1. Since the primary objective of the study was to assess user preferences in terms of satisfaction, no formal tutorial was provided. However, a moderator was available to offer additional explanations upon participant request. Furthermore, the information of agent models was not provided to participants.
D.2 Qualitative Analysis of Human Preference
Through the user study conducted in Section 5.1, we find diverse motivations underlying participants’ decisions. To gain a deeper understanding of user preferences, we conduct a manual categorization of participant feedback and count the frequency of each category. As shown in Figure 5, conciseness and successful task completion emerged as the most prominent factors contributing to user satisfaction. In addition, the coherence and naturalness of responses are also identified as significant influences. These qualitative findings are consistent with the quantitative results presented in Section 5.1, reinforcing the observation that while task completion performance plays a critical role, multiple other factors also contribute to overall user satisfaction.
Figure 5: Manual categorization of factors that affect user satisfaction during interactions.
Appendix E Additional Discussions on Agent-Mixing Experiments
$\tau$-Bench-Airline
[…]
mistral-large .383±0.038 .924±0.046 .581±0.022 .777±0.005 .358±0.029 .605
   + llama-3.3-70B interm. .452±0.032 .957±0.008 .562±0.021 .821±0.019 .238±0.048 .606
   + qwen2.5-72B interm. .417±0.031 .939±0.017 .545±0.029 .809±0.008 .336±0.050 .609
qwen2.5-72B .403±0.021 .951±0.015 .637±0.031 .821±0.008 .329±0.021 .628
   + llama-3.3-70B interm. .447±0.020 .958±0.014 .663±0.022 .831±0.008 .272±0.051 .634
   + mistral-large interm. .423±0.039 .933±0.032 .578±0.050 .776±0.009 .367±0.013 .615
llama-3.3-70B .463±0.055 .952±0.013 .643±0.035 .818±0.015 .263±0.056 .628
   + mistral-large interm. .446±0.054 .937±0.014 .457±0.071 .764±0.009 .333±0.032 .587
   + qwen2.5-72B interm. .404±0.033 .953±0.007 .599±0.016 .803±0.014 .361±0.019 .624
Best-of-N .469±0.019 .941±0.023 .612±0.045 .806±0.012 .334±0.038 .632
Table 7: Full results for mixing agent experiments in $\tau$-Bench-Airline. Agents indicated with ‘‘interm.’’ are utilized in an intermediate understanding component. The best performance of each configuration is bolded.
In this section, we discuss the potential feasibility of mixing agents according to the agentic pipeline in task planning scenarios. As mentioned in Table 4, we report the entire combination of three open-source models in Table 7.
Mixing with mistral-large
While mistral-large achieves higher task completion ($\mathcal{R}$) compared to llama-3.3-70B and qwen2.5-72B, it demonstrates substantially lower performance on PIPA metrics. Specifically, we observe a decline in task completion when substituting the intermediate understanding component with llama-3.3-70B (.358 $\rightarrow$ .238) or qwen2.5-72B (.358 $\rightarrow$ .336). Conversely, there is a marginal increase in the average PIPA scores (.605 $\rightarrow$ .606, .609);however, this slight improvement is unlikely to have a significant impact on overall user satisfaction.
Mixing with qwen2.5-72B
We find that qwen2.5-72B mixed with llama-3.3-70B for intermediate understanding yields the highest PIPA scores (.634) among all other rows, despite its relatively poor task completion rate (.272). Conversely, using mistral-large for intermediate understanding does not increase--and sometimes decreases--PIPA scores (.615), but it demonstrates the highest task completion scores (.367).
Mixing with llama-3.3-70B
As emphasized in Section 5.2, using qwen2.5-72B for intermediate understanding with llama-3.3-70B shows the highest improvement in task completion (.263 $\rightarrow$ .361) with only a minimal drop in the average PIPA score (.628 $\rightarrow$ .624). We observe similar, but not as impressive performance when we use mistral-large for intermediate understanding (task completion: .263 $\rightarrow$ .333; average PIPA: .628 $\rightarrow$ .587).
On top of these results, we conduct a supplementary experiment on mixing agents, referred to as the Best-of-N configuration in Table 7. In this setting, for each step (i.e., all states and responses), we collect the outputs of all three models and select the one that achieves the highest average performance across PIPA evaluation metrics (excluding task completion), as assessed by llama-3.3-70B judge. Although this configuration does not yield the highest performance in either task completion or the average PIPA score individually, it demonstrates a balanced performance between these two key indicators of user satisfaction.
Consequently, these findings reveal intriguing insights on model combinations, suggesting that there are lots more insights to uncover for combining the strengths of different models. We highlight a promising direction for future research.
Appendix F Erroneous Patterns in User Simulators
Proactivity / Goal-Seeking Errors
This describes when the user simulator does not demonstrate proactive and goal-seeking errors. More specifically, they do not follow up or confirm when the agent seems like their request was not explicitly addressed. For example:
  • When booking a flight, one of the requests from the user simulator was to book specific seats. The agent does not directly acknowledge this request, and when asking the user simulator to confirm the booking details, it does not mention anything about the seats. The user simulator should have followed up with the agent to confirm that their requests for the seat were also processed.
Instruction Contradiction
This describes instances where the user simulator directly contradicts an instruction given to them. Reasons for this include (1) explicit disregard for an instruction and (2) failure to recognize when conditional instructions apply to the current context. For example:
  • Instructions specified that the user simulator does not remember their reservation ID. Initially, the user simulator adhered and said that they do not remember it. However, when the agent said that they need it to proceed, the simulator claimed to remember it and started to provide made-up reservation IDs. This reveals poor alignment--both violating direct instructions and exhibiting deceptive behavior.
  • The user simulator was instructed to use a different form of payment when transaction costs exceeded \$100. However, when presented with this scenario, the simulator failed to recognize the applicability of the conditional instruction and proceeded with the wrong form of payment.
Missing Details
This describes when there is an attempt from the user simulator to follow an instruction, but it does so in the wrong format/order, omits key details, or forgets one or more instructions. For example:
  • The user instructions specified to mention multiple requirements at once and in a specific order, however, the user simulator brought them up independently or in different orders.
  • The user simulator was instructed to change the topic after 3 agent messages, but did it much later.
  • The user simulator was instructed to book a flight, and one of the instructions were to add extra carry-on baggage. After completing the flight booking as specified, the user simulator failed to ask to add the extra carry-on baggage despite the system asking if there was anything else they could help with.
Misinterpretation or Confusion
This is when the user simulator misreads or fails to interpret certain parts of the instruction. This can be, in part, due to providing unclear instructions. For example:
  • The user simulator instructions provide a name but do not explicitly specify that this is also their user ID. In some cases, the user simulator fails to recognize that their name is also their user ID.
Through our analysis in Section 5.3, we find that utilizing an LLM judge fails to identify numerous failure cases of the user simulator, necessitating reliance on human evaluators. However, such human-in-the-loop evaluations are both financially and logistically unsustainable. To mitigate this limitation, establishing a comprehensive and standardized protocol for a user simulator is essential (Aher et al., 2023; Yoon et al., 2024). We leave this to future work.